# Decomposing Global Feature Effects Based on Feature Interactions

### v2

@@ Abstract
Global feature effect methods, such as partial dependence plots, provide an intelligible visualization of the expected marginal feature effect. However, such global feature effect methods can be misleading, as they do not represent local feature effects of single observations well when feature interactions are present. We formally introduce generalized additive decomposition of global effects (GADGET), which is a new framework based on recursive partitioning to find interpretable regions in the feature space such that the interaction-related heterogeneity of local feature effects is minimized. We provide a mathematical foundation of the framework and show that it is applicable to the most popular methods to visualize marginal feature effects, namely partial dependence, accumulated local effects, and Shapley additive explanations (SHAP) dependence. Furthermore, we introduce and validate a new permutation-based interaction test to detect significant feature interactions that is applicable to any feature effect method that fits into our proposed framework. We empirically evaluate the theoretical characteristics of the proposed methods based on various feature effect methods in different experimental settings. Moreover, we apply our introduced methodology to three real-world examples to showcase their usefulness.
@@ 1 Introduction
Machine learning (ML) models are increasingly used in various application fields such as medicine (Shipp et al., 2002) or social sciences (Stachl et al., 2020). Their outstanding predictive performance comes from their ability to capture complex non-linear relationships and feature interactions in the data. However, this ability also complicates explaining the inner workings of an ML model. A lack of explainability might hurt trust or might even be a deal-breaker for high-stakes decisions (Lipton, 2018). Hence, ongoing research on model-agnostic interpretation methods to explain any ML model has grown quickly in recent years.
One promising type of explanation is produced by feature effect methods, which explain how features influence the model predictions similar to interpreting the coefficients of a linear model or visualizing the non-linear effects of a generalized additive model (GAM) through splines. We distinguish between local and global feature effect methods. Local feature effect methods—such as individual conditional expectation (ICE) curves (Goldstein et al., 2015) or Shapley values / Shapley additive explanations (SHAP) (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017)—explain how each feature influences the prediction of a single observation. In contrast, global feature effect methods explain the general model behavior. Since global feature effects of ML models are often non-linear, they are usually represented as marginal curves instead of single effect numbers; the most popular variants are partial dependence (PD) plots (Friedman, 2001), accumulated local effects (ALE) plots (Apley and Zhu, 2020), or SHAP dependence (SD) plots (Lundberg et al., 2019). These global curves are expressible via an aggregation of local counterparts (e.g., ICE curves are aggregated to a PD curve). However, such an aggregation might cause information loss due to heterogeneity in local feature effects (e.g., see Figure 1). This so-called aggregation bias is usually caused by feature interactions learned by the ML model, leading to a global feature effect that does not appropriately represent many individual observations in the data (Herbinger et al., 2022; Mehrabi et al., 2021). We refer to this heterogeneity as interaction-related heterogeneity. Consequently, global explanations can be misleading and cannot give a complete picture when feature interactions are present, e.g., see how the different shapes of ICE curves are not well represented by the global PD curve in Figure 1. This issue is particularly relevant when ML models are trained on biased data, as the model may, in turn, learn these biases (Mehrabi et al., 2021). Although local explanations could reveal the learned bias, it remains hidden in global explanations due to aggregation (e.g., see the COMPAS example in Section 8).
@@
We introduce GADGET, a novel framework, which provides regional explanations based on feature effects by partitioning the feature space into interpretable subspaces. We prove that GADGET minimizes feature interactions for a predefined set of features using any feature effect method satisfying the local decomposability axiom (Section 4.1 and 4.2). We show that popular feature effect methods (PD, ALE, and SD) satisfy the local decomposability axiom (Section 4.3-4.5) and illustrate their applicability within GADGET in Section 4.6 with instructions on how to estimate and visualize the corresponding regional feature effects in Appendix B.4. Moreover, we propose several measures to quantify feature interactions based on GADGET, which provide more insights into the learned effects and the remaining heterogeneity related to the interaction (Section 4.7). Furthermore, we introduce and validate the permutation interaction (PINT) test, the first statistical test to detect significant feature interactions in a model-agnostic fashion based on the underlying feature effect method (Section 5). We empirically evaluate the theoretical characteristics of the different methods based on several simulation settings (Sections 6 and 7) and illustrate their usefulness for two real-world examples in Section 8. We provide a third higher-dimensional example in Section 9 and also propose a general filtering procedure for such scenarios. All proposed methods and reproducible scripts for the experiments are available online via https://github.com/JuliaHerbinger/gadget/.
@@ 2.2 Definition of Interactions and the Functional ANOVA Decomposition
The functional ANOVA decomposition has already been studied by Stone (1994) to explain prediction models. It decomposes a square-integrable prediction function $\hat{f}(\mathbf{x})$ by
@@
[…]
@@ 2.3 Feature Effect Methods
The PD plot introduced by Friedman (2001) visualizes the marginal effect of a set of features $W\subseteq\{1,\ldots,p\}$ for a fitted model $\hat{f}$ by integrating over the joint distribution over all other features in $-W$, which we denote by $\mathbbm{P}(X_{-W})$. The number of features in $W$ is typically one or two. The theoretical PD function is defined by  [^3] The notation here is used consistently with Friedman and Popescu (2008). Please note: While the right-hand side of the definition always refers to the ML model $\hat{f}$, the $f$ in $f^{PD}_{W}$ does not refer to the true relationship (see Section 2.1); $f^{PD}_{W}$ is rather to be read as a new symbol. In the same manner, the “hat” of $\hat{f}^{PD}_{W}$ is not introduced to refer to $\hat{f}$ now – but expresses the fact that we now estimate based on data.
@@
which is estimated for $m$ grid points [^4] Instead of using all feature values $\mathbf{x}_{W}$, an equidistant grid or a grid based on randomly selected feature values or quantiles of $\mathbf{x}_{W}$ are commonly chosen (Molnar et al., 2022). denoted by $\tilde{\mathbf{x}}_{W}=\tilde{\mathbf{x}}_{W}^{(1)},\ldots,\tilde{\mathbf{x}}_{W}^{(m)}$ using Monte-Carlo integration: $\hat{f}_{W}^{PD}(\tilde{\mathbf{x}}_{W})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(\tilde{\mathbf{x}}_{W},\mathbf{x}^{(i)}_{-W})$. Hence, the PD plot visualizes the pairs $\{(\tilde{\mathbf{x}}_{W}^{(k)},\hat{f}_{W}^{PD}(\tilde{\mathbf{x}}_{W}^{(k)}))\}_{k=1}^{m}$. Goldstein et al. (2015) introduced ICE curves, which, for an observation $\mathbf{x}$, fixes the values $\mathbf{x}_{-W}$ to the observed ones, and varies the values $\mathbf{x}_{W}$ yielding the ICE function $\hat{f}^{\mathbf{x}}_{W}(\tilde{\mathbf{x}}_{W})=\hat{f}(\tilde{\mathbf{x}}_{W},\mathbf{x}_{-W})$, where the superscript $\mathbf{x}$ denotes the observation whose values $\mathbf{x}_{W}$ are replaced with $\tilde{\mathbf{x}}_{W}$. ICE plots visualize the pairs $\{(\tilde{\mathbf{x}}_{W}^{(k)},\hat{f}^{\mathbf{x}^{(i)}}_{W}(\tilde{\mathbf{x}}_{W})\}_{k=1}^{m}$, representing prediction changes of each observation as the values of features in $W$ vary. Notably, the PD curve is the point-wise average over all the ICE curves, i.e., $\hat{f}_{W}^{PD}(\tilde{\mathbf{x}}_{W})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}^{\mathbf{x}^{(i)}}_{W}(\tilde{\mathbf{x}}_{W})$. ICE curves can reveal heterogeneous effects caused by interactions between features $\mathbf{x}_{W}$ and other features, which remain hidden in the PD curve due to averaging (see Figure 1). However, they do not reveal which features interact with $\mathbf{x}_{W}$. In addition, if the model contains complex interactions, it becomes more difficult to identify clear patterns with interacting features (even if we color-code ICE curves as in Figure 1).
@@
The PD function integrates over the marginal distribution and is likely to extrapolate in empty or sparse regions of the feature space when features are correlated. This means that the prediction function is evaluated at unrealistic combinations of feature values in Eq. 2, which can lead to inaccurate PD estimates, especially for non-parametric models such as neural networks (Apley and Zhu, 2020). Hence, ALE plots are based on integrating over the conditional distribution to avoid extrapolation. The uncentered ALE ${f}^{ALE}_{W}(x)$ for $|W|=1$ at feature value $x_{W}\sim\mathds{P}(X_{W})$ and with $z_{0}=\min(\mathbf{x}_{W})$ is theoretically defined by
@@
Shapley values can be quantified using either a marginal-based (similar to PDs) or a conditional-based approach. While the former might encounter extrapolation issues in the presence of feature correlations, the latter presents the challenge of estimating the conditional distribution. For a more thorough discussion on this topic, see Appendix I.
@@ 2.4 Quantification of Interaction Effects
Visualization methods for feature effects are usually limited to a maximum of two features. To understand which feature effects depend on other features due to interactions, we are interested in detecting and quantifying the strength of feature interactions. Building on the mean-centered PD function $\hat{f}_{j}^{PD,c}(\tilde{\mathbf{x}}_{j})=\hat{f}_{j}^{PD}(\tilde{\mathbf{x}}_{j})-\frac{1}{m}\sum_{k=1}^{m}\hat{f}_{j}^{PD}(\tilde{x}_{j}^{(k)})$, Friedman and Popescu (2008) introduced the H-Statistic as a global interaction measure that quantifies the interaction strength between a feature of interest $\mathbf{x}_{j}$ and all other features $\mathbf{x}_{-j}$ by
@@ 3 Related Work
Regional explanations (Molnar et al., 2023; Britton, 2019; Hu et al., 2020; Scholbeck et al., 2024; Herbinger et al., 2022)—sometimes also referred to cohort explanations (Sokol and Flach, 2020)—explain subspaces of the entire feature space and thus bridge the gap between local and global explanations. In general, we can distinguish between unsupervised and supervised methods to define these subspaces. Concerning the former, Britton (2019) suggests the use of an unsupervised clustering approach to group ICE curves based on their partial derivatives and apply a decision tree post hoc for explanation purposes. A similar approach has been introduced by Zhang et al. (2021). However, Herbinger et al. (2022) showed that such unsupervised approaches can produce misleading interpretations. To overcome these downsides, other works such as Molnar et al. (2023); Scholbeck et al. (2024); Herbinger et al. (2022) propose methods that are based on a supervised decision tree approach to define regions that are interpretable and distinct. The first two references focus on finding regions with few feature correlations and few non-linearities, respectively. Instead, the REPID method (Herbinger et al., 2022) minimizes feature interactions within each region and decomposes the global PD into regional PDs, ensuring the homogeneity of ICE curves within each region. However, REPID has its limitations: First, REPID generates feature-specific regions by minimizing interactions between a particular feature and all other features. Applying the method to a different feature of interest will usually result in different regions, complicating interpretations when the effects of multiple features are of interest. Second, REPID relies on ICE and PD plots and thus inherits also their issues, such as extrapolation, potentially providing misleading interpretations if both feature interactions and correlations are present (Molnar et al., 2022). To overcome these limitations, we propose a more flexible framework in Section 4 that is capable of producing regions where interactions between multiple features are minimized and where different feature effect methods other than ICE and PD plots can be used.
Related Work on Significance Tests for Interactions.
@@
ANOVA is one of the most prominent statistical testing procedures to detect significant (two-way) interaction terms in statistical models. An algorithm to detect important feature interactions without a formal definition of a statistical test was introduced along with the standard functional ANOVA decomposition in Hooker (2004). Later, Mentch and Hooker (2017) introduced a statistical test for tree-based methods to detect additive structures and, thus, feature interactions. Henelius et al. (2016) introduced a statistical test for classification tasks to detect additive structures between different subsets of features based on their earlier work Henelius et al. (2014). In their approach, feature values of the considered feature subset are permuted jointly within each class to break possible interactions with other feature subsets while preserving interactions within the subset. However, this approach is limited to classification tasks and breaks correlations with features outside of the regarded subset. Consequently, learned spurious interactions are detected by the algorithm as true feature interactions. Furthermore, statistical tests for two-way interaction detection have been introduced by Loh (2002) or Sorokina et al. (2008). While the former provides a formal statistical test statistic, the latter is based on a heuristic to detect two-way interactions by comparing the performance of a full and restricted model. To the best of our knowledge, a model-agnostic statistical test designed for classification and regression tasks to detect interactions between a feature of interest $\mathbf{x}_{j}$ and the remaining features $\mathbf{x}_{-j}$, as presented in Section 5, has not yet been introduced.

### v3

@@ Abstract
Global feature effect methods, such as partial dependence plots, provide an intelligible visualization of the expected marginal feature effect. However, such global feature effect methods can be misleading, as they do not represent local feature effects of single observations well when feature interactions are present. We formally introduce generalized additive decomposition of global effects (GADGET), which is a new framework based on recursive partitioning to find interpretable regions in the feature space such that the interaction-related heterogeneity of local feature effects is minimized. We provide a mathematical foundation of the framework and show that it is applicable to the most popular methods to visualize marginal feature effects, namely partial dependence, accumulated local effects, and Shapley additive explanations (SHAP) dependence. Furthermore, we introduce and validate a new permutation-based interaction detection procedure that is applicable to any feature effect method that fits into our proposed framework. We empirically evaluate the theoretical characteristics of the proposed methods based on various feature effect methods in different experimental settings. Moreover, we apply our introduced methodology to three real-world examples to showcase their usefulness.
@@ 1 Introduction
Machine learning (ML) models are increasingly used in various application fields such as medicine (Shipp et al., 2002) or social sciences (Stachl et al., 2020). Their exceptional predictive performance stems from their ability to capture complex non-linear relationships and feature interactions in the data. However, this ability also complicates explaining the inner workings of an ML model. A lack of explainability might hurt trust or might even be a deal-breaker for high-stakes decisions (Lipton, 2018). Hence, ongoing research on model-agnostic interpretation methods to explain any ML model has grown quickly in recent years.
One promising type of explanation is produced by feature effect methods, which explain how features influence the model predictions similar to interpreting the coefficients of a linear model or visualizing the non-linear effects of a generalized additive model (GAM) through splines. We distinguish between local and global feature effect methods. Local feature effect methods—such as individual conditional expectation (ICE) curves (Goldstein et al., 2015) or Shapley values / Shapley additive explanations (SHAP) (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017)—explain how each feature influences the prediction of a single observation. In contrast, global feature effect methods explain the general model behavior. Since global feature effects of ML models are often non-linear, they are usually represented as marginal curves instead of single effect numbers; the most popular variants are partial dependence (PD) plots (Friedman, 2001), accumulated local effects (ALE) plots (Apley and Zhu, 2020), or SHAP dependence (SD) plots (Lundberg et al., 2020). These global curves are expressible via an aggregation of local counterparts (e.g., ICE curves are aggregated to a PD curve). However, such an aggregation might cause information loss due to heterogeneity in local feature effects (e.g., see Figure 1). This so-called aggregation bias is usually caused by feature interactions learned by the ML model, leading to a global feature effect that does not appropriately represent many individual observations in the data (Herbinger et al., 2022; Mehrabi et al., 2021). We refer to this heterogeneity as interaction-related heterogeneity. Consequently, global explanations can be misleading and cannot give a complete picture when feature interactions are present, e.g., see how the different shapes of ICE curves are not well represented by the global PD curve in Figure 1. This issue is particularly relevant when ML models are trained on biased data, as the model may, in turn, learn these biases (Mehrabi et al., 2021). Although local explanations could reveal the learned bias, it remains hidden in global explanations due to aggregation (e.g., see the COMPAS example in Section 8).
@@
We introduce GADGET, a novel framework, which provides regional explanations based on feature effects by partitioning the feature space into interpretable subspaces. We prove that GADGET minimizes feature interactions for a predefined set of features using any feature effect method satisfying the local decomposability axiom (Section 4.1 and 4.2). We show that popular feature effect methods (PD, ALE, and SD) satisfy the local decomposability axiom (Section 4.3-4.5) and illustrate their applicability within GADGET in Section 4.6 with instructions on how to estimate and visualize the corresponding regional feature effects in Appendix C.4. Moreover, we propose several measures to quantify feature interactions based on GADGET, which provide more insights into the learned effects and the remaining heterogeneity related to the interaction (Section 4.7). Furthermore, we introduce and validate the permutation-based interaction detection (PINT) procedure to detect feature interactions in a model-agnostic fashion based on the underlying feature effect method (Section 5). We empirically evaluate the theoretical characteristics of the different methods based on several simulation settings (Sections 6 and 7) and illustrate their usefulness for two real-world examples in Section 8. We provide a third higher-dimensional example in Section 9 and also propose a general filtering procedure for such scenarios. All proposed methods and reproducible scripts for the experiments are available online via https://github.com/JuliaHerbinger/gadget/.
@@ 2.2 Definition of Interactions and the Functional ANOVA Decomposition
This definition of feature interactions is closely related to the concept of functional ANOVA [^1] This should not be confused with the classical ANOVA (Analysis of Variance), which is a statistical method that tests for significant differences among group means in a population. decomposition, which has already been studied by Sobol’ (1993); Stone (1994) to explain black box functions. The functional ANOVA decomposition aims to decompose a square-integrable prediction function $\hat{f}(\mathbf{x})$ into a sum of lower-order functions:
@@
The standard functional ANOVA decomposition by Hooker (2004) assumes independent features and ensures that the component functions $g_{W}(\mathbf{x}_{W})$ can be optimally and uniquely defined by imposing the so-called vanishing condition [^2] See Appendix A for more details. (Li and Rabitz, 2012; Rahman, 2014). This decomposition is widely used in sensitivity analysis to further uniquely decompose the variance of each component from Eq. (1) and derive a feature importance measure (see Sobol indices in Sobol’, 1993; Owen, 2013). Hooker (2007) introduced the generalized functional ANOVA decomposition with a relaxed vanishing condition\@footnotemark that leads to a unique decomposition in case of dependent features. These conditions guarantee orthogonality between the components representing different feature subsets and ensure that each component is distinct without redundancy. In this paper, we do not attempt to compute the exact decomposition presented in Eq. (1); instead, we rely on the existence of such a decomposition for our theoretical proofs in Section 4. We refer to Appendix A for more details on the functional ANOVA decomposition and related research regarding its estimation.
@@ 2.3 Feature Effect Methods
The PD plot introduced by Friedman (2001) visualizes the marginal effect of a set of features $W\subseteq\{1,\ldots,p\}$ for a fitted model $\hat{f}$ by integrating over the joint distribution over all other features in $-W$. The number of features in $W$ is typically one or two. The theoretical PD function is defined by  [^3] The notation here is used consistently with Friedman and Popescu (2008). Please note: While the right-hand side of the definition always refers to the ML model $\hat{f}$, the $f$ in $f^{PD}_{W}$ does not refer to the true relationship (see Section 2.1); $f^{PD}_{W}$ is rather to be read as a new symbol. In the same manner, the “hat” of $\hat{f}^{PD}_{W}$ is not introduced to refer to $\hat{f}$ now – but expresses the fact that we now estimate based on data.
@@
which is estimated using Monte-Carlo integration: $\hat{f}_{W}^{PD}(\tilde{\mathbf{x}}_{W})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(\tilde{\mathbf{x}}_{W},\mathbf{x}^{(i)}_{-W})$ at a specific grid point $\tilde{\mathbf{x}}_{W}$. For constructing a PD curve, we typically use $m$ grid points [^4] Instead of using all observed feature values $\mathbf{x}_{W}$, an equidistant grid or a grid based on randomly selected feature values or quantiles of $\mathbf{x}_{W}$ are commonly chosen (Molnar et al., 2022). denoted by $\tilde{\mathbf{x}}_{W}^{(1)},\ldots,\tilde{\mathbf{x}}_{W}^{(m)}$ and visualize the pairs $\{(\tilde{\mathbf{x}}_{W}^{(k)},\hat{f}_{W}^{PD}(\tilde{\mathbf{x}}_{W}^{(k)}))\}_{k=1}^{m}$. Goldstein et al. (2015) introduced ICE curves, which, for an observation $\mathbf{x}$, fixes the values $\mathbf{x}_{-W}$ to the observed ones, and varies the values $\mathbf{x}_{W}$ yielding the ICE function $\hat{f}^{\mathbf{x}}_{W}(\tilde{\mathbf{x}}_{W})=\hat{f}(\tilde{\mathbf{x}}_{W},\mathbf{x}_{-W})$, where the superscript $\mathbf{x}$ denotes the observation whose values $\mathbf{x}_{W}$ are replaced with $\tilde{\mathbf{x}}_{W}$. ICE plots visualize the pairs $\{(\tilde{\mathbf{x}}_{W}^{(k)},\hat{f}^{\mathbf{x}^{(i)}}_{W}(\tilde{\mathbf{x}}_{W})\}_{k=1}^{m}$, representing prediction changes of each observation as the values of features in $W$ vary. Notably, the PD curve is the point-wise average over all the ICE curves, i.e., $\hat{f}_{W}^{PD}(\tilde{\mathbf{x}}_{W})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}^{\mathbf{x}^{(i)}}_{W}(\tilde{\mathbf{x}}_{W})$. ICE curves can reveal heterogeneous effects caused by interactions between features $\mathbf{x}_{W}$ and other features, which remain hidden in the PD curve due to averaging (see Figure 1). However, they do not reveal which features interact with $\mathbf{x}_{W}$. In addition, if the model contains complex interactions, it becomes more difficult to identify clear patterns with interacting features (even if we color-code ICE curves as in Figure 1).
@@
The PD function uses the marginal distribution and is likely to extrapolate in empty or sparse regions of the feature space when features are correlated. This means that the prediction function is evaluated at unrealistic combinations of feature values in Eq. (2), which can lead to inaccurate PD estimates, especially for models with high flexibility in capturing complex feature interactions, such as neural networks (Apley and Zhu, 2020). Hence, ALE uses the conditional distribution to avoid extrapolation. The uncentered ALE ${f}^{ALE}_{W}(x)$ for $|W|=1$ at feature value $x_{W}\sim\mathds{P}_{X_{W}}$ and with $z_{0}=\min(\mathbf{x}_{W})$ is defined by
[…]
@@
Shapley values can be quantified using either a marginal-based (similar to PDs) or a conditional-based approach. While the former might encounter extrapolation issues in the presence of feature correlations, the latter presents the challenge of estimating the conditional distribution. For a more thorough discussion on this topic, see Appendix J.
@@ 2.4 Quantification of Interaction Effects
Visualization methods for feature effects are usually limited to a maximum of two features. To understand which feature effects depend on other features due to interactions, we are interested in detecting and quantifying the strength of feature interactions. Building on the mean-centered PD function $\hat{f}_{j}^{PD,c}(\tilde{x}_{j})=\hat{f}_{j}^{PD}(\tilde{x}_{j})-\frac{1}{m}\sum_{k=1}^{m}\hat{f}_{j}^{PD}(\tilde{x}_{j}^{(k)})$ at each grid point $\tilde{x}_{j}\in\{\tilde{x}_{j}^{(1)},\dots,\tilde{x}_{j}^{(m)}\}$, Friedman and Popescu (2008) introduced the H-Statistic as a global interaction measure that quantifies the interaction strength between a feature of interest $\mathbf{x}_{j}$ and all other features $\mathbf{x}_{-j}$ by
@@ 3 Related Work
Regional explanations (Molnar et al., 2023; Britton, 2019; Hu et al., 2020; Scholbeck et al., 2024; Herbinger et al., 2022)—sometimes also referred to cohort explanations (Sokol and Flach, 2020)—explain subspaces of the entire feature space, thereby bridging the gap between local and global explanations. In general, we can distinguish between unsupervised and supervised methods to define these subspaces. Concerning the former, Britton (2019) suggests the use of an unsupervised clustering approach to group ICE curves based on their partial derivatives and apply a decision tree post hoc for explanation purposes. A similar approach has been introduced by Zhang et al. (2021). However, Herbinger et al. (2022) showed that such unsupervised approaches can produce misleading interpretations. To overcome these downsides, other works such as Molnar et al. (2023); Scholbeck et al. (2024); Herbinger et al. (2022) propose methods that are based on a supervised decision tree approach to define regions that are interpretable and distinct. The first two references focus on finding regions with few feature correlations and few non-linearities, respectively. Instead, the REPID method (Herbinger et al., 2022) minimizes feature interactions within each region and decomposes the global PD into regional PDs, ensuring the homogeneity of ICE curves within each region. However, REPID has its limitations: First, REPID generates feature-specific regions by minimizing interactions between a particular feature and all other features. Applying the method to a different feature of interest typically results in different regions, complicating interpretations when multiple features are of interest. Second, REPID relies on ICE and PD plots and thus inherits also their issues, such as extrapolation, potentially providing misleading interpretations if both feature interactions and correlations are present (Molnar et al., 2022). To overcome these limitations, we propose a more flexible framework in Section 4 that is capable of producing regions where interactions between multiple features are minimized and where different feature effect methods other than ICE and PD plots can be used.
Related Work on Detecting Feature Interactions.
@@
ANOVA is one of the most prominent statistical testing procedures to detect significant (two-way) interaction terms in statistical models. An algorithm to detect important feature interactions without a formal definition of a statistical test was introduced along with the standard functional ANOVA decomposition in Hooker (2004). Later, Mentch and Hooker (2017) introduced a statistical test for tree-based methods to detect additive structures and, thus, feature interactions. Henelius et al. (2016) introduced a statistical test for classification tasks to detect additive structures between different subsets of features based on their earlier work Henelius et al. (2014). In their approach, feature values of the considered feature subset are permuted jointly within each class to break possible interactions with other feature subsets while preserving interactions within the subset. However, this approach is limited to classification tasks and breaks correlations with features outside of the regarded subset. Consequently, learned spurious interactions are detected by the algorithm as true feature interactions. Furthermore, statistical tests for two-way interaction detection have been introduced by Loh (2002) or Sorokina et al. (2008). While the former provides a formal statistical test statistic, the latter is based on a heuristic to detect two-way interactions by comparing the performance of a full and restricted model. To the best of our knowledge, a model-agnostic procedure designed for classification and regression tasks to detect interactions between a feature of interest $\mathbf{x}_{j}$ and the remaining features $\mathbf{x}_{-j}$, as presented in Section 5, has not yet been introduced.